OpenStudy (anonymous):

Evaluate. The upper limit of an antiderivative is 2 and the lower limit is 1. the integral is 5*sqrt(x)*ln(5)/2*sqrt(x)

OpenStudy (anonymous):

do you mean the integrand?

OpenStudy (mimi_x3):

Is it: \[\large \int\limits\frac{5\sqrt{x}*\ln5}{2\sqrt{x}} \] ?

OpenStudy (anonymous):

even I am getting confused. Post the question properly.

OpenStudy (wasiqss):

mimi i guess yes this is the form

OpenStudy (mimi_x3):

But it doesn't look right..

OpenStudy (wasiqss):

professor we need to work on this person's typo too :P

OpenStudy (mimi_x3):

Hehe yeah; for him/her to overcome it just like you. :P

OpenStudy (anonymous):

if mimi is right, then sqrt x will get cancelled and it will become the easiest integration ever! :P

OpenStudy (wasiqss):

yes akshat we know that :P

OpenStudy (anonymous):

lol..wat if he lanad8 meant \[5\sqrt{x} [\ln 5/2] \sqrt{x} \] ???

OpenStudy (anonymous):

even thats easy buy at least we can integrate x! lol :P

OpenStudy (mimi_x3):

Yeah..we have to wait for lanad8 to respond!

OpenStudy (anonymous):

sorry for the confusion. thanks, zakaullah

OpenStudy (anonymous):

WElcome man